OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____. (13/34, -15/34) (3/22, -7/22) (13/22, 1/22)

OpenStudy (anonymous):

\[\huge{7x-3y=4}\] \[\huge{2x-4y=1}\] 1) \[\huge{x=\frac{4+3y}{7}}\] 2) \[\huge{2(\frac{{4+3y}}{7})-4y=1}\] \[\huge{\frac{8+6y-28y}{7}=1}\] \[\huge{\frac{8-22y}{7}=1}\] \[\huge{8-22y=7}\] \[\huge{y=\frac{1}{22}}\]

OpenStudy (anonymous):

so \[x =4+3\left( 1/22 \right) \over7\]

OpenStudy (anonymous):

= 13/22